# Consistency Models Made Easy

### v1

@@ Abstract
Consistency models (CMs) are an emerging class of generative models that offer faster sampling than traditional diffusion models. CMs enforce that all points along a sampling trajectory are mapped to the same initial point. But this target leads to resource-intensive training: for example, as of 2024, training a SoTA CM on CIFAR-10 takes one week on 8 GPUs. In this work, we propose an alternative scheme for training CMs, vastly improving the efficiency of building such models. Specifically, by expressing CM trajectories via a particular differential equation, we argue that diffusion models can be viewed as a special case of CMs with a specific discretization. We can thus fine-tune a consistency model starting from a pre-trained diffusion model and progressively approximate the full consistency condition to stronger degrees over the training process. Our resulting method, which we term Easy Consistency Tuning (ECT), achieves vastly improved training times while indeed improving upon the quality of previous methods: for example, ECT achieves a 2-step FID of 2.73 on CIFAR10 within 1 hour on a single A100 GPU, matching Consistency Distillation trained of hundreds of GPU hours. Owing to this computational efficiency, we investigate the scaling law of CMs under ECT, showing that they seem to obey classic power law scaling, hinting at their ability to improve efficiency and performance at larger scales. Code is available.
@@ 1 Introduction
Diffusion Models (DMs) [21, 63], or Score-based Generative Models (SGMs) [65, 66], have vastly changed the landscape of visual content generation with applications in images [56, 57, 23, 11, 17, 55], videos [6, 4, 2, 22, 15], and 3D objects [54, 70, 36, 8, 1]. DMs progressively transform a data distribution to a known prior distribution (e.g. Gaussian noise) according to a stochastic differential equation (SDE) [66] and train a model to denoise noisy observations. Samples can be generated via a reverse-time SDE that starts from noise and uses the trained model to progressively denoise it. However, sampling from a DM naively requires hundreds to thousands of model evaluations due to the curvature of the diffusion sampling trajectory [26], making the entire generative process slow. Many approaches have been proposed to address this issue, including training-based techniques such as distillation [44, 14, 58, 60, 74, 50, 13], adaptive compute architectures for the backbone model [49, 69], as well as training-free methods such as fast samplers [34, 42, 75, 80] or interleaving small and large backbone models during sampling [52]. However, the speedup achieved by these sampling techniques usually comes at the expense of the quality of generated samples.
@@
Consistency Models (CMs) [67, 64] are a new family of generative models, closely related to diffusion models, that have demonstrated promising results as faster generative models. These models learn a mapping between noise and data that all the points of the sampling trajectory map to the same initial data point. Owing to this condition, these models are capable of generating high-quality samples in 1-2 model evaluations. The best such models so far, built using improved Consistency Training (iCT) [64], have pushed the quality of images generated by 1-step CMs trained from scratch to a level comparable with SoTA DMs using thousands of steps for generation. Unfortunately, CMs remain time-consuming and practically challenging to train: the best practice so far takes many times longer than similar-quality diffusion models while involving complex hyperparameter choices in the training process. In total, this has substantially limited the uptake of CMs within the community.
In this work, we propose a vastly more efficient scheme for building consistency models. Specifically, we first rewrite the consistency condition, through the lens of dynamical systems, as a finite-difference approximation of a particular differential equation, which we call the differential consistency condition. Directly training consistency models using this condition as a loss function raises several challenges; we propose several strategies (including choosing a specific discretization curriculum and using a particular weighting function) to address them. Most importantly, using this formalism, we observe that DMs can be seen as a special case of CMs with a loose form of differential consistency condition. This lets us smoothly interpolate from DM to CM by progressively tightening the consistency condition, gradually making the time discretization smaller as training progresses. Combining these strategies results in a streamlined training recipe, called Easy Consistency Tuning (ECT), that lets us build a consistency model starting from a pretrained diffusion model, and thereby significantly decreasing the training cost.
ECT achieves better 2-step sample quality on CIFAR-10 [35] and ImageNet 64$\times$64 [10] than the prior art, while using less than $\mathbf{2\%}$ of the training FLOPs; even if we factor in the pre-training budget of the corresponding diffusion model, ECT only requires $\nicefrac{{1}}{{4}}$ the time of the state-of-the-art iCT [64]. ECT also significantly reduces the inference cost by $\nicefrac{{1}}{{1000}}$ compared to Pretrained Score SDE/DMs [66] while maintaining comparable generation quality.
@@ 2 Preliminaries
Let $p_{\text{data}}({\mathbf{x}}_{0})$ denote the data distribution. Diffusion models (DMs) perturb this distribution by adding monotonically increasing i.i.d. Gaussian noise with standard deviation $\sigma(t)$ from $t=0$ to $T$ such that $p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})={\mathcal{N}}({\mathbf{x}}_{0},\sigma^{2}(t){\bm{I}})$, and $\sigma(t)$ is chosen such that $\sigma(0)=\sigma_{\text{min}}$ and $\sigma(T)=\sigma_{\text{max}}$. This process is described by the following SDE [66]
@@
where $\mathbf{w}$ is the standard Wiener process, $\mathbf{f}(\cdot,t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, and $g(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient. Samples can be generated by solving the reverse-time SDE starting from $t=T$ to 00 and sampling ${\mathbf{x}}_{T}\sim{\mathcal{N}}(0,\sigma^{2}_{\text{max}}{\bm{I}})$. Song et al. [66] show that this SDE has a corresponding ODE, called the probability flow ODE (PF-ODE), whose trajectories share the same marginal probability densities as the SDE. We follow the notation in Karras et al. [26] to describe the ODE as
@@
where $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ denotes the score function. Prior works [26, 67] set $\sigma(t)=t$ which gives
@@
Note that the consistency function needs to satisfy the boundary condition at $t=0$. Prior works [67, 64, 26] impose this boundary condition by parametrizing the CM as
@@
where $\theta$ is the model parameter, $F_{\theta}$ is the network to train, and $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are time-dependent scaling factors such that $c_{\text{skip}}(0)=1$, $c_{\text{out}}(0)=0$. This parameterization guarantees the boundary condition by design. We discuss specific choices of $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ in Appendix C.
@@
In the seminal work, Song et al. [67] use Learned Perceptual Similarity Score (LPIPS) [76] as a metric function, set $w(t_{i})=1$ for all $t_{i}$, and sample $t_{i}$ according to the sampling scheduler by  Karras et al. [26]: $t_{i}=\left(t_{\text{max}}^{1/\rho}+\frac{i}{N-1}(t_{\text{min}}^{1/\rho}-t_{\text{max}}^{1/\rho})\right)^{\rho}$ for $i\in{\mathcal{U}}[1,N-1]$ and $\rho=0.7$. Further, the score function $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t})$ can either be estimated from a pretrained diffusion model, which results in Consistency Distillation (CD), or can be estimated with an unbiased score estimator
@@
The follow-up work, iCT [64], introduces several improvements that significantly improve training efficiency as well as the performance of CMs. First, the LPIPS metric, which introduces undesirable bias in generative modeling, is replaced with a Pseudo-Huber metric. Second, the network $f_{\theta^{-}}$ does not maintain an EMA of the past values of $f_{\theta}$. Third, iCT replaces the uniform weighting scheme $w(t_{i})=1$ with $w(t_{i})=\frac{1}{t_{i+1}-t_{i}}$. Further, the scaling factors of noise embeddings and dropout are carefully selected. Fourth, iCT introduces a complex discretization curriculum during training:
@@
3 Easy Consistency Tuning (ECT)
We will first introduce the differential consistency condition that forms the core of our method to train CMs and derive a loss objective based on this condition. Next, we will analyze this loss objective and highlight some challenges of training CMs with it. Based on this analysis, we present our method, Easy Consistency Tuning (ECT). ECT is a simple, principled approach to efficiently train CMs to meet the (differential) consistency condition. The resulting CMs can generate high-quality samples in 1 or 2 sampling steps.
@@ 3.1 Differential Consistency Condition
As previously stated in Sec. 2, CMs learn a consistency function $f({\mathbf{x}}_{t},t)$ that maps the noisy image ${\mathbf{x}}_{t}$ back to the clean image ${\mathbf{x}}_{0}$: $f({\mathbf{x}}_{t},t)={\mathbf{x}}_{0}$. Instead, we argue that CMs can be defined through the time derivative of the model output, given by the differential form
@@
Since the clean image ${\mathbf{x}}_{0}$ is independent of the noise level $t$, its time derivative is zero.
However, the differential form $\frac{\mathrm{d}f}{\mathrm{d}t}=0$ alone is not sufficient to guarantee that the model output will match the clean image, as there exist trivial solutions where the model maps all the inputs to a constant value, such as $f({\mathbf{x}}_{t},t)\equiv 0$. To eliminate these collapsed solutions, we follow Song et al. [67], Song and Dhariwal [64] and impose a boundary condition for $f({\mathbf{x}}_{t},t)={\mathbf{x}}_{0}$:
@@
To satisfy the consistency condition, we minimize the distance between $f_{\theta}({\mathbf{x}}_{t})$ and $f_{\theta}({\mathbf{x}}_{r})$. For the boundary condition ($t>r=0$), we optimize $f_{\theta}({\mathbf{x}}_{r})$ to align with the clean image ${\mathbf{x}}_{0}$. For higher noise levels ($t>r>0$), we freeze $f_{\operatorname{sg}(\theta)}({\mathbf{x}}_{r})$ using the stop-gradient operator $\operatorname{sg}$ and optimize $f_{\theta}({\mathbf{x}}_{t})$ to align with $f_{\operatorname{sg}(\theta)}({\mathbf{x}}_{r})$.
@@
where we can extract a weighting function $w(t,r)=\frac{1}{t-r}$, ${\mathbf{x}}_{t}={\mathbf{x}}_{0}+t\cdot\bm{\epsilon}$, ${\mathbf{x}}_{r}={\mathbf{x}}_{0}+r\cdot\bm{\epsilon}$ using a shared noise direction $\bm{\epsilon}\sim p(\bm{\epsilon})$, and ${\mathbf{d}}(\cdot,\cdot)$ is a metric function. In practice, we set ${\mathbf{d}}(\cdot,\cdot)$ to the squared $L_{2}$ metric. Further, we discuss an interesting relation between the Pseudo-Huber metric proposed in Song and Dhariwal [64] and the weighting function in  Sec. 3.3. To regularize the model, dropout masks should also be consistent across $f_{\theta}({\mathbf{x}}_{t})$ and $f_{\theta}({\mathbf{x}}_{r})$ by setting a shared random seed, ensuring that the noise levels $t$ and $r$ are the only varying factors.
@@ 3.2 The "Curse of Consistency" and its Implications
The objective in Eq. 12 can be challenging to optimize when $\Delta t\to 0$. This is because the prediction errors from each discretization interval accumulate, leading to slow training convergence or, in the worst case, divergence. To further elaborate, consider a large noise level $T$. We first follow the notation of CT [67] and split the noise horizon $[0,T]$ into $N$ smaller consecutive subintervals. We can error bound $1$-step prediction of CM as
@@
where $r_{1}=0<t_{1}=r_{2}<\cdots<t_{N-2}=r_{N}<t_{N}=T$, and $e_{\max}=\max_{i}\|f_{\theta}({\mathbf{x}}_{t_{i}})-f_{\theta}({\mathbf{x}}_{r_{i}})\|$, for $i=1,\cdots,N$. Ideally, we want both $N$ and $e_{\max}$ to be small so that this upper bound is small, but in practice, there is a trade-off between these two terms. As $\Delta t_{i}=(t_{i}-r_{i})\to 0$, $e_{\max}$ decreases because ${\mathbf{x}}_{t_{i}}$ and ${\mathbf{x}}_{r_{i}}$ will be close, and it is easier to predict both $f_{\theta}({\mathbf{x}}_{r_{i}})$ and $f_{\theta}({\mathbf{x}}_{t_{i}})$. However, $N$ will increase as $\Delta t_{i}\to 0$. In contrast, for a large $\Delta t_{i}$, vice-versa holds true. It is difficult to theoretically estimate the rate at which $e_{\max}$ decreases when $\Delta t_{i}=t_{i}-r_{i}\to 0$, as it depends on the optimization process—specifically, how effectively we train the model to minimize the consistency error in each interval. If $e_{\max}$ decreases more slowly than $N$, the product of the two can increase instead, resulting in a worse prediction error $\|f_{\theta}({\mathbf{x}}_{T})-{\mathbf{x}}_{0}\|$.
@@
The key insight from the above observation is that if we train from scratch, strictly following the differential form $\nicefrac{{\mathrm{d}f}}{{\mathrm{d}t}}=0$ using a tiny $\Delta t_{i}\approx 0$, the resulting model can get stuck during training due to the accumulated consistency errors from each interval $\Delta t_{i}$. We verify this hypothesis by training a series of CMs using different numbers of intervals $N$, and corresponding $\Delta t$ of consistency condition as shown in Fig. 1.
3.3 Tackling the "Curse of Consistency"
The discussion in Sec. 3.2 highlighted the training instability issue that might arise when we train CMs from scratch with the loss objective in Eq. 12 while directly following $\Delta t\approx 0$. However, these issues can be mitigated with the right design choices. In this section, we list several strategies that address the aforementioned training instability issues and help improve the efficiency of CMs.
Start ECT with a pretrained diffusion model.
Drawing inspiration from iCT [64], we start ECT with a large $\Delta t$, and gradually shrink $\Delta t\to 0$. In our problem setup, Since $t>r\geq 0$, we have the largest possible $\Delta t=t$ with $r=0$, which yields

### v2

@@ Abstract
Consistency models (CMs) offer faster sampling than traditional diffusion models, but their training is resource-intensive. For example, as of 2024, training a state-of-the-art CM on CIFAR-10 takes one week on 8 GPUs. In this work, we propose an effective scheme for training CMs that largely improves the efficiency of building such models. Specifically, by expressing CM trajectories via a particular differential equation, we argue that diffusion models can be viewed as a special case of CMs. We can thus fine-tune a consistency model starting from a pretrained diffusion model and progressively approximate the full consistency condition to stronger degrees over the training process. Our resulting method, which we term Easy Consistency Tuning (ECT), achieves vastly reduced training times while improving upon the quality of previous methods: for example, ECT achieves a 2-step FID of 2.73 on CIFAR10 within 1 hour on a single A100 GPU, matching Consistency Distillation trained for hundreds of GPU hours. Owing to this computational efficiency, we investigate the scaling laws of CMs under ECT, showing that they obey the classic power law scaling, hinting at their ability to improve efficiency and performance at larger scales. Our code is publicly available, making CMs more accessible to the broader community.
@@ 1 Introduction
Diffusion Models (DMs) (Ho et al., 2020, Song et al., 2021a), or Score-based Generative Models (SGMs) (Song et al., 2020; 2021b), have vastly changed the landscape of visual content generation with applications in images (Rombach et al., 2021, Saharia et al., 2022, Ho et al., 2022a, Dhariwal and Nichol, 2021, Hatamizadeh et al., 2023, Ramesh et al., 2021), videos (Brooks et al., 2024, Blattmann et al., 2023, Bar-Tal et al., 2024, Ho et al., 2022b, Gupta et al., 2023), and 3D objects (Poole et al., 2022, Wang et al., 2024a, Lee et al., 2024, Chen et al., 2024, Babu et al., 2023). DMs progressively transform a data distribution to a known prior distribution (e.g. Gaussian noise) according to a stochastic differential equation (SDE) (Song et al., 2021b) and train a model to denoise noisy observations. Samples can be generated via a reverse-time SDE that starts from noise and uses the trained model to progressively denoise it. However, sampling from a DM naively requires hundreds to thousands of model evaluations due to the curvature of the diffusion sampling trajectory (Karras et al., 2022), making the entire generative process slow. Many approaches have been proposed to address this issue, including training-based techniques such as distillation (Luhman and Luhman, 2021, Salimans and Ho, 2022, Luo et al., 2024, Gu et al., 2023, Sauer et al., 2023, Geng et al., 2024, Yin et al., 2023, Nguyen and Tran, 2023), adaptive compute architectures for the backbone model (Moon et al., 2023, Tang et al., 2023), as well as training-free methods such as fast samplers (Kong and Ping, 2021, Lu et al., 2022a, Zhang and Chen, 2022, Zhou et al., 2023, Xue et al., 2024) or interleaving small and large backbone models during sampling (Pan et al., 2024). However, the speedup achieved by these sampling techniques usually comes at the expense of the quality of generated samples.
@@
Consistency Models (CMs) (Song et al., 2023) are a new family of generative models, closely related to diffusion models, that have demonstrated promising results as faster generative models. These models learn a mapping between noise and data, and all the points of the sampling trajectory map to the same initial data point. Owing to this condition, consistency models are capable of generating high-quality samples in 1-2 model evaluations. The best such models so far, built using improved Consistency Training (iCT) (Song and Dhariwal, 2023), have pushed the quality of images generated by 1-step CMs trained from scratch to a level comparable with SoTA DMs using thousands of steps for sampling. Unfortunately, CMs remain time-consuming and practically challenging to train: the best practice takes many times longer than similar-quality DMs while involving complex hyperparameter choices in the training process. In total, this has substantially limited the uptake of CMs within the community.
In this work, we introduce a differential perspective on consistency models, leading to the formulation of the differential consistency condition in continuous time. This insight reveals the link between diffusion models and consistency models, viewing DMs as a special case of CMs with loose discretization. This observation motivates us to smoothly interpolate from DM to CM by progressively tightening the consistency condition, bootstrapping pretrained DMs to 1-step CMs w/o using extra frozen teachers. We term this strategy as Easy Consistency Tuning (ECT), which includes diffusion pretraining as a special stage of the continuous time training schedule.
ECT significantly improves both training efficiency and performance. On ImageNet 64$\times$64 (Deng et al., 2009), ECT achieves superior 1-step and 2-step sample quality compared to the prior art. Similarly, on CIFAR-10 (Krizhevsky, 2009) 2-step sample quality of ECT surpasses previous methods. The total cost of the pretraining and tuning scheme requires only $\nicefrac{{1}}{{4}}\sim\nicefrac{{1}}{{3}}$ of the computational resources (FLOPs) used by the current state-of-the-art method, iCT (Song and Dhariwal, 2023), while the tuning stage can be remarkably lightweight, typically accounting for $10\%$ or less of the overall cost and further benefiting from scaling.
Leveraging ECT’s computational efficiency, we conduct the first study into the scaling behaviors of CMs, revealing the classic power law scaling for model size, FLOPs, and training compute. The scaling also suggests a sweet spot of using smaller few-step CMs over larger 1-step CMs in certain scenarios. This computational efficiency enables us to explore the design space of CMs by using the tuning stage as a proxy. Notably, tuning findings, such as weighting functions, can improve the pretraining stage and, in turn, enhance the overall pretraining + tuning pipeline for CMs.
In short, we summarize our contributions as follows:
  • We develop Easy Consistency Tuning (ECT), a pretraining + tuning scheme for training CMs in continuous time, demonstrating significant efficiency and performance gains compared to the current best practices for training CMs.
  • We investigate CMs’ scaling behaviors for the first time and reveal the classic power law.
  • We explore the design space of CMs through ECT, introducing the continuous-time schedule and better weighting functions for CMs.
@@ 2 Preliminaries
Let $p_{\text{data}}({\mathbf{x}}_{0})$ denote the data distribution. Diffusion models (DMs) perturb this distribution by adding monotonically increasing i.i.d. Gaussian noise with standard deviation $\sigma(t)$ from $t=0$ to $T$ such that $p_{t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})={\mathcal{N}}({\mathbf{x}}_{0},\sigma^{2}(t){\bm{I}})$, and $\sigma(t)$ is chosen such that $\sigma(0)=\sigma_{\text{min}}$ and $\sigma(T)=\sigma_{\text{max}}$. This process is described by the following SDE (Song et al., 2021b)
@@
where $\mathbf{w}$ is the standard Wiener process, $\mathbf{f}(\cdot,t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, and $g(\cdot):\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient. Samples can be generated by solving the reverse-time SDE starting from $t=T$ to 00 and sampling ${\mathbf{x}}_{T}\sim{\mathcal{N}}(0,\sigma^{2}_{\text{max}}{\bm{I}})$. (Song et al., 2021b) show that this SDE has a corresponding ODE, called the probability flow ODE (PF-ODE), whose trajectories share the same marginal probability densities as the SDE. We follow the notation in (Karras et al., 2022) to describe the ODE as
@@
where $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ denotes the score function. Prior works (Karras et al., 2022, Song et al., 2023) set $\sigma(t)=t$ which yields
@@
Note that the consistency function needs to satisfy the boundary condition at $t=0$. Prior works (Song et al., 2023, Song and Dhariwal, 2023, Karras et al., 2022) impose this boundary condition by parametrizing the CM as
@@
where $\theta$ is the model parameter, $F_{\theta}$ is the network to train, and $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are time-dependent scaling factors such that $c_{\text{skip}}(0)=1$, $c_{\text{out}}(0)=0$. This parameterization guarantees the boundary condition by design. We discuss specific choices of $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ in Appendix D.
@@
In the seminal work, (Song et al., 2023) use Learned Perceptual Similarity Score (LPIPS) (Zhang et al., 2018) as a metric function, set $w(t_{i})=1$ for all $t_{i}$, and sample $t_{i}$ according to the sampling scheduler by  (Karras et al., 2022): $t_{i}=\left(t_{\text{max}}^{1/\rho}+\frac{i}{N-1}(t_{\text{min}}^{1/\rho}-t_{\text{max}}^{1/\rho})\right)^{\rho}$ for $i\in{\mathcal{U}}[1,N-1]$ and $\rho=0.7$. Further, the score function $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t})$ can either be estimated from a pretrained diffusion model, which results in Consistency Distillation (CD), or can be estimated with an unbiased score estimator
@@
The follow-up work, iCT (Song and Dhariwal, 2023), introduces several improvements that significantly improve training efficiency as well as the performance of CMs. First, the LPIPS metric, which introduces undesirable bias in generative modeling, is replaced with a Pseudo-Huber metric. Second, the network $f_{\theta^{-}}$ does not maintain an EMA of the past values of $f_{\theta}$. Third, iCT replaces the uniform weighting scheme $w(t_{i})=1$ with $w(t_{i})=\frac{1}{t_{i+1}-t_{i}}$. Further, the scaling factors of noise embeddings and dropout are carefully selected. Fourth, iCT introduces a complex discretization curriculum during training:
@@
3 Probing consistency models
We will first introduce the differential consistency condition and the loss objective based on this condition. Next, we will analyze this loss objective and highlight the challenges of training CMs with it. Based on this analysis, we present our method, Easy Consistency Tuning (ECT). ECT is a simple, principled approach to efficiently train CMs to meet the (differential) consistency condition. The resulting CMs can generate high-quality samples in 1 or 2 sampling steps.
@@ 3.1 Differential Consistency Condition
As stated in Sec. 2, CMs learn a consistency function $f({\mathbf{x}}_{t},t)$ that maps the noisy image ${\mathbf{x}}_{t}$ back to the clean image ${\mathbf{x}}_{0}$: $f({\mathbf{x}}_{t},t)={\mathbf{x}}_{0}$. Instead, by taking the time derivative of both sides, given by the differential form
@@
However, the differential form $\frac{\mathrm{d}f}{\mathrm{d}t}=0$ alone is not sufficient to guarantee that the model output will match the clean image, as there exist trivial solutions where the model maps all the inputs to a constant value, such as $f({\mathbf{x}}_{t},t)\equiv 0$. To eliminate these collapsed solutions, (Song et al., 2023, Song and Dhariwal, 2023) impose the boundary condition for $f({\mathbf{x}}_{t},t)={\mathbf{x}}_{0}$ via model parameterization:
@@
To satisfy the consistency condition, we minimize the distance between $f_{\theta}({\mathbf{x}}_{t})$ and $f_{\theta}({\mathbf{x}}_{r})$. For the boundary condition, we optimize $f_{\theta}({\mathbf{x}}_{t})$ to align with the clean image ${\mathbf{x}}_{0}$. For noise levels $t>r>0$, we freeze $f_{\operatorname{sg}(\theta)}({\mathbf{x}}_{r})$ using the stop-gradient operator $\operatorname{sg}$ and optimize $f_{\theta}({\mathbf{x}}_{t})$ on higher variance inputs to align with self teacher’s prediction $f_{\operatorname{sg}(\theta)}({\mathbf{x}}_{r})$ on lower variance inputs (Chen and He, 2021).
@@
where we can extract a weighting function $w(t,r)=\frac{1}{t-r}$, ${\mathbf{x}}_{t}={\mathbf{x}}_{0}+t\cdot\bm{\epsilon}$, ${\mathbf{x}}_{r}={\mathbf{x}}_{0}+r\cdot\bm{\epsilon}$ using a shared noise direction $\bm{\epsilon}\sim p(\bm{\epsilon})$, and ${\mathbf{d}}(\cdot,\cdot)$ is a metric function. This formulation generalizes Song et al. (2023)’s discrete training schedule to continuous time. To regularize the model, dropout masks should also be consistent across $f_{\theta}({\mathbf{x}}_{t})$ and $f_{\theta}({\mathbf{x}}_{r})$ by setting a shared random seed, ensuring that the noise levels $t$ and $r$ are the only varying factors.
@@ 3.2 The "Curse of Consistency" and its Implications
The objective in Eq. 12 can be challenging to optimize when $\Delta t\to 0$. This is because the prediction errors from each discretization interval accumulate, leading to slow training convergence or, in the worst case, divergence. To further elaborate, consider a large noise level $T$. We first follow the notation of CT (Song et al., 2023) and split the noise horizon $[0,T]$ into $N$ smaller consecutive subintervals. We can error bound the error of $1$-step prediction as
@@
where $r_{1}=0<t_{1}=r_{2}<\cdots<t_{N-2}=r_{N}<t_{N}=T$, and $e_{\max}=\max_{i}\|f_{\theta}({\mathbf{x}}_{t_{i}})-f_{\theta}({\mathbf{x}}_{r_{i}})\|$, for $i=1,\cdots,N$. Ideally, we want both $N$ and $e_{\max}$ to be small so that this upper bound is small. But in practice, there is a trade-off between these two terms. As $\Delta t_{i}=(t_{i}-r_{i})\to 0$, $e_{\max}$ decreases because ${\mathbf{x}}_{t_{i}}$ and ${\mathbf{x}}_{r_{i}}$ will be close, and it is easier to predict both $f_{\theta}({\mathbf{x}}_{r_{i}})$ and $f_{\theta}({\mathbf{x}}_{t_{i}})$. However, $N$ will increase as $\Delta t_{i}\to 0$. In contrast, for a large $\Delta t_{i}$, vice-versa holds true. It is difficult to theoretically estimate the rate at which $e_{\max}$ decreases when $\Delta t_{i}=t_{i}-r_{i}\to 0$, as it depends on the optimization process—specifically, how effectively we train the model to minimize the consistency error in each interval. If $e_{\max}$ decreases more slowly than $N$, the product of the two can increase instead, resulting in a worse prediction error $\|f_{\theta}({\mathbf{x}}_{T})-{\mathbf{x}}_{0}\|$.
The insight from the above observation is that when training from scratch by strictly following the differential form $\nicefrac{{\mathrm{d}f}}{{\mathrm{d}t}}=0$ with a tiny $\Delta t_{i}\approx 0$, the resulting model might converge slowly due to accumulated consistency errors from each interval $\Delta t_{i}$. To investigate this hypothesis, we conducted experiments training a series of CMs using different fixed numbers of intervals $N$, and corresponding $\Delta t$ values for the consistency condition, as illustrated in Fig. 1. Our observations align with Song and Dhariwal (2023), with the key distinction that we use a fixed $N$ for this analysis, as this approach provides better insight on how much the precise approximation of the differential consistency condition matters and isolates the effect of discretization errors.
3.3 Easy Consistency Tuning (ECT)
The discussion in Sec. 3.2 highlighted the training instability issue that can arise when we naively train CMs from scratch and directly optimize for the differential consistency condition with the loss objective in Eq. 12 while directly following $\Delta t\approx 0$. In this section, we propose several strategies that largely alleviate the aforementioned training issues and improve the efficiency of CMs. We term this approach Easy Consistency Tuning (ECT), as it effectively balances training stability and model performance while simplifying the CM training process. ECT follows a two-stage approach: diffusion pretraining, followed by consistency tuning, which we will detail in the following subsections.
Diffusion Pretraining + Consistency Tuning.
Drawing inspiration from iCT’s adaptive discrete-time schedule (Song and Dhariwal, 2023), we start ECT with a large $\Delta t$, and gradually shrink $\Delta t\to 0$. In our problem setup, Since $t>r\geq 0$, we have the largest possible $\Delta t=t$ with $r=0$, which yields